OpenStudy (anonymous):

10. You want to buy a house in 9 years. The estimated cost is $175,800.00. You want to make a 20% down payment and closing costs are 7%. How much in total do you need save each month to reach your goal if you cover closing costs and the down payment? The closing costs are messing me up a bit any help?

OpenStudy (anonymous):

So closing costs are extra costs that you incur Closing costs are 7% of the cost of the house Closing costs = .07 *175800

OpenStudy (anonymous):

.07*175,800= 12,306*.20= 2,461.20 is what you need to save each month?

OpenStudy (anonymous):

Ok so there are 2 diff costs that you will have to save up for: #1 The closing cost which is .07*175,800 #2 The down payment which is .20*175,800

OpenStudy (anonymous):

So his goal is to save up money for 9 years so that he can cover the closing costs and the down payment

OpenStudy (anonymous):

closing cost is 12,306 and down payment is 35,160?

OpenStudy (anonymous):

Yes so the total cost he needs to save up is 12306+35160=?

OpenStudy (anonymous):

both added are 47,466... how do i get the monthy savings for those 9 years?

OpenStudy (anonymous):

Ok so there are 12 months in a year and he saves up for 9 years So the total amount of months in 9 years are 12*9=108 r Do you follow?

OpenStudy (anonymous):

He needs to accumulate 47,466 in 9 years or rather he needs to accumulate 47466 in 108 months

OpenStudy (anonymous):

so... 47,466 divided by 108?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so 439.50 a month? thanks!

OpenStudy (anonymous):

yea that sounds right